### v1

@@ 4.3.2 Performance by Machine Difficulties
To evaluate the model’s performance across different complexity levels, we clustered the scores based on machine difficulties. This allows us to identify the areas where the model performs better or worse depending on the complexity. Table  10 shows that our model CIPHER performs better on average due to its distinguished high scores on easy and insane machines, with 56.25 and 50.96 points respectively, creating the highest gaps of about 1.69 and 2.22 points to the second-place model. Llama 3 70B takes second place, with slightly better scores on medium and hard machines by 0.52 and 0.43 points. This demonstrates our model’s strength in penetration testing reasoning guidance compared to other penetration testing models, even those with ten times the parameter size.
Table 10: Model Performance FARR Flow Reasoning Test Sorted by Machine Difficulty Scores
@@
whiterabbitneo-7b 33.75 30.01 28.47 28.75 30.25
lily-7b 40.73 34.66 34.50 32.03 35.48
qwen-7b 44.68 41.31 42.02 41.88 42.47
colibri-0.1-8b 49.74 47.38 45.82 48.75 47.92
hermes2pro-llama3-8b 50.43 47.76 48.54 48.37 48.78
cipher-chaco-1.5-dpo-7b 52.80 47.96 46.43 47.69 48.72
mistral-ins-0.2-7b 51.21 49.34 47.27 47.89 48.93
qwen1.5-72b 52.02 47.81 50.88 47.67 49.60
gpt-3.5-turbo 53.41 48.18 49.67 47.72 49.74
pentestai-7b 52.82 48.52 49.97 48.70 50.00
openhermes-2.5-7b 53.05 48.51 50.27 48.50 50.08
llama-3-70b-ins 54.56 51.28 51.48 48.74 51.52
CIPHER DPO 7B 56.25 50.80 51.05 50.96 52.26
@@ 4.4 PurpleLlama CyberSecEval
We also evaluate our model using the most popular cyber security evaluation for LLM, PurpleLlama CybersecEval2 Bhatt et al. (2024). This evaluation was initially designed to measure the safety of the LLM model before being released. Since we are developing the model that accurately suggests the offensive exploitation technique, we can also measure how well our model is capable of attacks listed in MITRE ATT&CK. The value that we show in Figure 10 is the malicious score, unlike the original paper that showcases the benign score.
Figure 10: MITRE ATT&ACK Capabilities Result
@@
On the contrary, we observe that the performance of general models tends to peak in reconnaissance and discovery techniques, even outperforming CIPHER in these areas. This is because reconnaissance and service discovery are common topics that are not considered dangerous. However, using only high-quality augmented write-ups and raw context-hacking techniques material, CIPHER enhances the MITRE ATT&CK capabilities of its base model and base conversation dataset. While this benchmark does not represent the model’s full performance, it showcases its extensive knowledge coverage across different attack techniques.
@@ 5.1 Limitation
Since CIPHER is designed to understand the current situation and suggest subsequent actions, we did not heavily emphasize coding datasets. CIPHER’s coding performance can be enhanced by changing the base model to a coding model or using a higher parameter base model. From a data-centric perspective, improving coding troubleshooting performance by creating synthetic datasets is possible. This can be done similarly to the CIPHER method, involving experts to fix code and provide explanations, reasoning, and examples.
Our FARR Flow Augmentation is imperfect; some generated augmentations still mention "the author" when they are supposed to be, just like finding notes. We have explicitly forbidden this behavior, but it is still in some samples. This could be caused by the limitation of the model used for generation. However, this does not affect the model response when augmented into question, even if it exists, it will only cause minor issues in perspective sentences.
Another difficulty in FARR Flow augmentation is providing challenging "Findings." Generated findings often contain hints, such as "found a zip file, need to crack the password." While this challenges the model to determine which tools to use for cracking, it already suggests a line of reasoning. As a result, the model’s focus will be on cracking the password instead of first checking the metadata or trying known target user passwords. This nuance needs to be understood by the model generating the FARR Flow. Although this is not a direct limitation of FARR Flow augmentation, it opens up future work to refine the pipeline and eliminate such results. Rather than waiting for AGI-level LLM, a feasible approach is to use agents or more advanced chain algorithms.
In the FARR Flow evaluation, we use Llama 3 70B as the judge LLM. Even though it is a state-of-the-art model, it has limitations and biases. Different judges can introduce different biases. Using multiple-judge LLMs in future work is a possible solution to reduce this bias.
@@ 5.2 Scaling
The first strategy involves scaling the dataset. This can be improved by incorporating more diverse and specific augmentations for the same dataset, using techniques like Orca Mukherjee et al. (2023) and Evolve instruct Xu et al. (2023), allowing the model to generalize better. The current cost of dataset augmentation ranges from \$800 to \$1000, excluding experimental costs, which can reach around \$10,000 for full augmentation.
@@ 5.4 Ethical Consideration
6 Conclusion
Our research aimed at developing an AI assistant chatbot capable of accurately guiding beginner penetration testers through the pentesting process. We utilized synthetic novice and expert pentesting conversations, raw text from high-quality documentation, and open-source general task datasets. The synthetic conversations were meticulously crafted with selected components to mimic expert responses to user queries. Additionally, we introduced a novel augmentation process called FARR Augmentation to generate high-quality information about penetration testing processes using write-ups as input (FARR Flow).
We designed synthetic questions to trigger the model’s understanding of the user’s situation and to measure its ability to suggest accurate actions. Performance was compared between our model, CIPHER, and other open-source penetration testing models, as well as higher parameter models. Our findings indicate that CIPHER slightly decreases general capabilities but significantly improves penetration testing abilities.
CIPHER provides more accurate and helpful suggestions compared to other penetration testing models across all machine question categories. For instance, CIPHER performs better than Llama 3 70B in Easy and Insane machine categories, likely because it excels in understanding both extreme cases and common penetration testing scenarios. However, it slightly underperforms in the Medium and Hard machine categories.
Future improvements for CIPHER could involve utilizing full dataset augmentation and employing larger parameter models. Furthermore, the FARR Flow method could be enhanced by crafting more complex questions or developing a virtual vulnerable machine LLM evaluation based on the FARR Flow steps.
@@ 6 Conclusion
Conceptualization, D.P.; methodology, D.P.; software, D.P., N.S. and A.A.A.; validation, D.P., N.S., A.A.A., T.T.H.L., I.M., A.Y.K.; formal analysis, D.P., N.S., A.A.A., T.T.H.L., A.Y.K.; investigation, D.P.; resources, D.P.; data curation, D.P.; writing—original draft preparation, D.P.; writing—review and editing, D.P., T.T.H.L., A.A.A, N.S., A.Y.K.; visualization, D.P., T.T.H.L., and A.A.A; supervision, T.T.H.L. and H.K.; project administration, H.K.; funding acquisition, H.K. All authors have read and agreed to the published version of the manuscript.
@@
This research was supported by the MSIT(Ministry of Science and ICT), Korea, under the Convergence security core talent training business(Pusan National University) support program(IITP-2024-2022-0-01201) supervised by the IITP(Institute for Information & Communications Technology Planning & Evaluation).
@@
Not applicable

### v2

@@ 4.3.2 Performance by Machine Difficulties
To evaluate the model’s performance across different complexity levels, we clustered the scores based on machine difficulties. This allows us to identify the areas where the model performs better or worse depending on the complexity. Table 10 shows that our model CIPHER performs better on average due to its distinguished high scores on easy and insane machines, with 56.25 and 50.96 points, respectively, creating the highest gaps of about 1.69 and 2.22 points from the second-place model. Llama 3 70B takes second place, with slightly better scores on medium and hard machines by 0.52 and 0.43 points. This demonstrates our model’s strength in penetration testing reasoning guidance compared to other penetration testing models, even those with ten times the parameter size.
In the detailed results for easy-difficulty machines shown in Figure 13, the accuracy scores of various models on easy-difficulty Hack The Box vulnerable machines are compared. CIPHER dominates, outperforming Llama 3 70B with an accuracy of 56.25. This is attributed to CIPHER’s extensive training with vulnerable machines in its dataset, which enhances its ability to generalize in these easier challenges, typically involving single-step problem-solving or a single vulnerability.
Table 10: Model performance FARR Flow reasoning test sorted by machine difficulty scores.
@@
whiterabbitneo-7b Whi 33.75 30.01 28.47 28.75 30.25
lily-7b seg 40.73 34.66 34.50 32.03 35.48
qwen1.5-7b Qwe 44.68 41.31 42.02 41.88 42.47
colibri-0.1-8b Cyb 49.74 47.38 45.82 48.75 47.92
hermes2pro-llama3-8b Nou 50.43 47.76 48.54 48.37 48.78
mistral-ins-0.2-7b Mis 51.21 49.34 47.27 47.89 48.93
qwen1.5-72b Qwe 52.02 47.81 50.88 47.67 49.60
gpt-3.5-turbo Ouyang et al. (2022) 53.41 48.18 49.67 47.72 49.74
openhermes-2.5-7b Teknium (2023) 53.05 48.51 50.27 48.50 50.08
llama-3-70b-ins Met 54.56 51.28 51.48 48.74 51.52
CIPHER DPO 7B (Ours) 56.25 50.80 51.05 50.96 52.26
{adjustwidth}
[…]
Figure 13: Easy machine performance results.
Following this scenario, in the medium machine challenge shown in Figure 14, CIPHER experienced a slight performance drop compared to Llama 3 70B. However, despite this small difference, CIPHER still achieved the highest score, creating a notable gap compared to models with 7B parameters. This suggests that CIPHER provides better guidance than the smaller models. It is important to note that Llama 3 70B’s high score primarily serves as a benchmark to gauge CIPHER’s competitive ability against larger parameter models. Additionally, since Llama 3 70B is used as the evaluation model, its own score may be influenced by bias.
This phenomenon is also observed in the hard machine challenge in Figure 15, where CIPHER achieves the best score among the 7B models but still falls below the average performance of Llama 3 70B. We believe that, in addition to potential bias, there may be inherent limitations in the 7B model’s ability to generalize reasoning and provide accurate guidance. These limitations could potentially be mitigated through dataset scaling and increasing the model size.
Moving on to the insane difficulty machine’s case in Figure 16, multiple pieces of information must be combined and a chained exploitation plan is required. Surprisingly, CIPHER leads the benchmark with a 2-point gap over all other models. This demonstrates CIPHER’s significant penetration testing guidance capabilities in extremely challenging scenarios that demand out-of-the-box exploitation, where even higher-parameter models reach their limits.
{adjustwidth}
[…]
Figure 14: Medium machine performance results.
{adjustwidth}
[…]
Figure 15: Hard machine performance results.
@@
Overall, CIPHER achieved the highest total average score, demonstrating superior performance in both the easy and insane tasks. It also excelled in the medium and hard tasks compared to other pentesting models of its size, placing second only to Llama 3 70B. This may be due to the fact that Llama 3 70B was also the evaluator, and its larger parameter size could have contributed to its advantage. Additionally, we discuss potential improvements for CIPHER in Section 5.2.
Figure 16: Insane machine performance results.
@@ 4.4 PurpleLlama CyberSecEval
We also evaluate our model using the most popular cyber security evaluation for LLMs, PurpleLlama CybersecEval2 Bhatt et al. (2024). This evaluation was initially designed to measure the safety of the LLM model before being released. Since we are developing a model that accurately suggests offensive exploitation techniques, we can also measure how well our model is capable of attacks listed in MITRE ATT&CK. The value that we show in Figure 17 is the malicious score, unlike the original paper that showcases the benign score.
@@
In contrast, we observe that the performance of general models tends to peak in reconnaissance and discovery techniques, even outperforming CIPHER in these areas. This is because reconnaissance and service discovery are common topics that are not considered dangerous. However, using only high-quality augmented write-ups and raw context-hacking techniques material, CIPHER enhances the MITRE ATT&CK capabilities of its base model and base conversation dataset. While this benchmark does not represent the model’s full performance, it showcases its extensive knowledge coverage across different attack techniques.
4.5 Real-World Beginner Use Case Experiment
In this experiment, we explore and analyze how a beginner penetration tester interacts with CIPHER to carry out common tasks such as reconnaissance, vulnerability discovery, and privilege escalation. The goal of this experiment is to evaluate the effectiveness of CIPHER in guiding new penetration testers through the initial stages of a penetration test, while also assessing how well the LLM adapts to novice-level queries and provides appropriate responses. Even though CIPHER has demonstrated strong performance in FARR Flow, this experiment aims to further validate its practical usability in real-world penetration testing scenarios, ensuring that it can handle practical questions effectively.
4.5.1 Reconnaissance
We begin the experiment with a typical beginner’s interaction with CIPHER during a reconnaissance phase. The user is prompted to ask for guidance on the first step in assessing the target system’s open ports. Figure 18 demonstrates CIPHER’s suggestion of using Nmap, a popular network scanning tool, to enumerate open ports on the target system. The model provides the appropriate command syntax nmap -p- --min-rate and also explains its relevance to the overall penetration testing process, emphasizing the need to identify all active services and open ports quickly for further exploitation without saving to any file.
The next scenario is involved with domain fuzzing, as shown in Figure 19. Providing an actionable guide for using ffuf in subdomain fuzzing, CIPHER effectively breaks down each component of the command. It explains common parameters like -u for the target URL, -H for adding custom headers, and -w for specifying wordlists, showcasing a solid understanding of both the tool and its practical applications. The use of placeholders such as FUZZ in the Host header is a standard fuzzing technique, and the inclusion of -fw 18 and -mc all shows a keen awareness of optimizing results and capturing a wide range of HTTP responses.
The response is highly effective for both novice and intermediate users, balancing simplicity with technical detail. CIPHER’s guidance is structured well, leaving room for customization, and demonstrates proficiency in subdomain enumeration. It could be enhanced with tips on handling challenges like rate-limiting and optimizing wordlists, but overall, it successfully conveys core principles and best practices as a penetration testing assistant.
In the next scenario, the user reports discovering port 5985 open on the target system and asks for advice on utilizing a long list of usernames and passwords. Figure 20 shows CIPHER’s response, providing a concise explanation of port 5985 being associated with Windows Remote Management (WinRM). CIPHER recommends using crackmapexec to automate testing combinations of usernames and passwords against the WinRM service, guiding the user with clear command syntax: crackmapexec winrm <target-ip> -d <domain-name> -u <username-list> -p <password-list> and explaining each element’s role in the process.
Figure 18: Nmap suggestion to scan all ports.
Figure 19: FFUF subdomain fuzzing.
Figure 20: CrackMapExec bruteforcing WinRM service.
This guidance is technically sound and user-friendly, offering both a practical tool for automating brute force attempts and explaining the relevance of port 5985. The response includes all the essential details, such as why this method is suitable for WinRM, making it a valuable resource to learn penetration testing. The combination of clear instructions and technical insight highlights CIPHER’s effectiveness as an educational tool in this context.
4.5.2 Tools Suggestion
In this scenario, the user inquires about the appropriate tool for connecting to WinRM on a target system. Figure 21 shows CIPHER’s response, recommending the use of Evil-WinRM, a tool specifically designed for penetration testing. CIPHER explains that Evil-WinRM provides a reliable shell for remote command execution and offers clear command syntax: evil-winrm -i <target-ip> -u <username> -p <password>, where the user inputs the target IP, username, and password.
Figure 21: Evil-WinRM as tool suggestion.
This interaction demonstrates CIPHER’s deep understanding of penetration testing techniques. The model not only suggests the correct tool but also explains its relevance, emphasizing the need for valid credentials when working with WinRM-enabled systems. By providing clear instructions, CIPHER ensures that even novice users can confidently execute the task, making it a valuable educational resource for penetration testers.
In this scenario, the user discovers that both FTP and HTTP ports are open on the target, with the ability to upload files anonymously to the FTP server. Figure 22 shows CIPHER’s response, which acknowledges the progress and suggests a common strategy: uploading a webshell. This is a script that allows command execution on the server via a web interface. CIPHER advises the user to first check the server’s configuration to determine which file extensions are allowed for uploads, as this will dictate the type of webshell that can be used. For example, if the server supports PHP, the user could upload a PHP webshell to gain control over the system.
Figure 22: FTP webshell recommendation.
CIPHER’s guidance shows an understanding of the penetration testing process, particularly when dealing with misconfigured FTP and HTTP services. The suggestion to check file extension permissions reflects the nuanced approach needed when attempting to escalate access, ensuring that the webshell method is tailored to the target’s configuration.
In the next case study, the user asks how to generate a reverse shell for a 64-bit Windows target in a DLL format. Figure 23 illustrates CIPHER’s detailed response, recommending the use of msfvenom from the Metasploit framework. CIPHER begins by advising the user to check the system’s architecture using the systeminfo command, ensuring the correct architecture set is used. Once confirmed, CIPHER provides a command syntax to generate the DLL reverse shell. The user is instructed to replace placeholders with appropriate values to create the reverse shell, which can then be used in the penetration testing process.
Figure 23: Exploit creation suggestion.
This response reflects a solid understanding of reverse shell generation, covering both technical accuracy and practical guidance. CIPHER explains not only how to generate the reverse shell but also emphasizes verifying the target’s architecture, demonstrating a nuanced approach to penetration testing and ensuring that the process runs smoothly.
4.5.3 Attack Vector Suggestion
In this scenario, the user asks about potential attack vectors for a website that accepts URLs and converts the content into PDF format. Figure 24 displays CIPHER’s response, which suggests several potential attack vectors, starting with Server-Side Request Forgery (SSRF). This attack allows the attacker to manipulate the server into fetching resources from unintended locations, possibly exposing internal services or sensitive data. CIPHER also points out the security risks associated with the PDF generation process, which could be exploited via Cross-Site Scripting (XSS) or SQL Injection if not properly secured. The model emphasizes the importance of gathering as much information as possible about the target system through enumeration before launching any attacks, demonstrating a nuanced understanding of web-based vulnerabilities.
Figure 24: Attack vector suggestion.
CIPHER’s response in Figure 25 provides a clear and practical guide to gaining code execution via WordPress, offering step-by-step instructions on file manipulation, such as editing the ‘404.php’ file and uploading a theme. The attention to detail ensures that users of all experience levels can follow along easily, with the model demonstrating a solid understanding of WordPress-specific vulnerabilities and how themes can be exploited for remote code execution.
Figure 25: Wordpress RCE instruction.
The choice to use a 404.php file and inject the system(\$_REQUEST[’cmd’]); payload shows smart tactics, as this file is often overlooked. While the response could enhance stealth advice, such as clearing logs or minimizing site disruption, it effectively balances clarity and technical depth, making it suitable for both beginners and more advanced users. CIPHER’s output showcases a well-rounded understanding of penetration testing within WordPress environments.
4.5.4 Privilege Escalation Suggestion
Shifting to the next example response in Figure 26, CIPHER offers a comprehensive, systematic approach to privilege escalation after gaining a foothold as a local user, aligning with established penetration testing methodologies. Starting with enumeration using tools like ‘ps’, ‘netstat’, ‘id’, and ‘sudo -l’, CIPHER ensures the user collects key system information, laying a strong foundation for identifying privilege escalation opportunities. The explanation of SUID/SGID files and how to locate them with ‘find’ is particularly valuable, providing actionable advice for both novice and experienced users.
CIPHER further demonstrates its depth by covering misconfigurations, kernel exploits, and shell escape techniques, which are crucial for exploiting system weaknesses. The model’s attention to advanced scenarios, such as credential reuse and inspecting insecure cron jobs, along with its focus on maintaining stealth throughout the process, shows a strong understanding of real-world penetration testing tactics. This balance of clarity and technical depth makes CIPHER’s guidance highly effective for beginners.
The next study in Figure 27 showcases CIPHER’s solid understanding of privilege escalation by helping a user exploiting Git’s ‘PAGER‘ environment variable to execute arbitrary shell commands. The command sudo PAGER=’sh -c ‘‘exec sh 0<&1’’’ git -p help takes advantage of Git’s behavior to spawn a shell, effectively bypassing restricted environments. CIPHER explains each component of the command clearly, ensuring that even users unfamiliar with environment variables can follow along, while demonstrating technical depth and practicality.
Figure 26: Privilege escalation advice after foothold first gained.
Figure 27: Privilige escalation using misconfigured Git.
The response balances usability and insight, making it highly accessible to beginners. CIPHER’s guidance is clear, actionable, and well-aligned with established privilege escalation techniques. A minor enhancement could include addressing potential limitations in more restrictive environments, but overall, the quality of CIPHER’s output is excellent, effectively supporting penetration testers to gain highest privilege.
Based on this experiment, we have proven that CIPHER can be effectively used as a practical penetration testing assistant, guiding users through the utilization of multiple tools from reconnaissance to post-exploitation. We observed that while CIPHER performs well overall, there are areas for improvement, such as providing more detailed explanations, example cases, or alternative commands. These improvements can be achieved by scaling up the dataset to support longer and more detailed responses. There is also available room for improvement in model size; despite this, CIPHER is able to respond accurately and assist users effectively.
@@ 5.1 Limitations
Since CIPHER is designed to understand the current situation and suggest subsequent actions, we did not heavily emphasize coding datasets. CIPHER’s coding performance can be enhanced by changing the base model to a coding model or using a higher-parameter base model. From a data-centric perspective, improving coding troubleshooting performance by creating synthetic datasets is possible. This can be achieved similarly to the CIPHER method, involving experts to fix code and provide explanations, reasoning, and examples.
Our FARR Flow augmentation is imperfect; some generated augmentations still mention “the author” when they are supposed to be just like finding notes. We have explicitly forbidden this behavior, but it is still in some samples. This could be caused by the limitation of the model used for generation. However, this does not affect the model response when augmented into a question, and even if it exists, it will only cause minor issues in perspective sentences.
Another difficulty in FARR Flow augmentation is providing challenging findings. Generated findings often contain hints, such as “found a zip file, need to crack the password”. While this challenges the model to determine which tools to use for cracking, it already suggests a line of reasoning. As a result, the model’s focus will be on cracking the password instead of first checking the metadata or trying known target user passwords. This nuance needs to be understood by the model generating the FARR Flow. Although this is not a direct limitation of FARR Flow augmentation, it opens up future work to refine the pipeline and eliminate such results. Rather than waiting for AGI-level LLM, a feasible approach is to use agents or more advanced chain algorithms.
In the FARR Flow evaluation, we use Llama 3 70B as the judge LLM. Even though it is a state-of-the-art model, it has limitations and biases. Different judges can introduce different biases. Using multiple judge LLMs in future work is a possible solution to reduce this bias.
@@ 5.2 Scaling
The first strategy involves scaling the dataset. This can be improved by incorporating more diverse and specific augmentations for the same dataset, using techniques like Orca Mukherjee et al. (2023) and Evolve instruct Xu et al. (2023), allowing the model to generalize better. Also, more contextual details can be added into the dataset for the model to understand details of the current chunk to be generated such as date, scope of the context, or even source metadata, preventing the model hallucinating on unknown context. In addition, wide coverage of penetration testing knowledge can be obtained by adding CVE, CWE, and CAPEC as resources for CIPHER dataset augmentation. Currently, the cost of CIPHER dataset augmentation ranges from USD 800 to USD 1000, excluding experimental costs, which can reach around USD 10,000 for full augmentation with the current dataset coverage.
@@ 5.3 Measuring Other Penetration Testing Capabilities
For future work, we plan to deeply explore how CIPHER can assist in uncontrolled environments, where system behaviors are unpredictable or information is incomplete. This exploration will allow us to assess CIPHER’s ability to adapt to irregular scenarios, such as misconfigured services or unexpected vulnerabilities. By comparing its performance in both controlled and uncontrolled environments, we aim to evaluate how well CIPHER can identify alternative actions and adapt its predictions when faced with less predictable conditions. This investigation will provide valuable insights into the robustness and potential of CIPHER’s integration into FARR Flow’s next-step prediction accuracy testing.
@@ 5.4 Ethical Consideration
6 Conclusions
Our research focused on developing CIPHER, an AI chatbot designed to guide beginner penetration testers through the pentesting process. We utilized synthetic novice and expert conversations, high-quality documentation, and open-source task datasets. Synthetic questions were crafted to assess the model’s understanding and ability to suggest accurate actions.
Our novel FARR Flow evaluation revealed that CIPHER slightly decreases general capabilities but significantly enhances pentesting skills, providing more accurate suggestions than other models on average, particularly in the easy and insane machine categories. However, it shows slight underperformance in the medium and hard categories, but still leads among the 7B parameter models.
In practical testing, CIPHER demonstrated it can assist effectively with its current parameter size, handling a wide range of tasks from reconnaissance to post-exploitation. There is room for improvement, particularly in providing more detailed explanations and alternative actions.
Future enhancements could involve full dataset augmentation, additional security knowledge base, and larger parameter models, as well as refining the FARR Flow method by crafting more complex questions or developing a virtual vulnerable machine LLM evaluation based on FARR Flow steps.
@@ 6 Conclusions
Conceptualization, D.P.; methodology, D.P.; software, D.P., N.S. and A.A.A.; validation, D.P., N.S., A.A.A., T.-T.-H.L., M.I. and A.Y.K.; formal analysis, D.P., N.S., A.A.A., T.-T.-H.L. and A.Y.K.; investigation, D.P.; resources, D.P.; data curation, D.P.; writing—original draft preparation, D.P.; writing—review and editing, D.P., T.-T.-H.L., A.A.A., N.S. and A.Y.K.; visualization, D.P., T.-T.-H.L. and A.A.A.; supervision, T.-T.-H.L. and H.K.; project administration, H.K.; funding acquisition, H.K. All authors have read and agreed to the published version of the manuscript.
@@
This research was supported by the MSIT (Ministry of Science and ICT), Korea, under the Convergence Security Core Talent Training Business (Pusan National University) support program (No. RS-2022-II221201) and was supported by the ITRC (Information Technology Research Center) support program (IITP-2024-RS-2020-II201797), both supervised by the IITP (Institute for Information & Communications Technology Planning & Evaluation).
@@
Not applicable.
[…]
Not applicable.
[…]
The data and benchmark presented in this study are available publicly at the our repository (https://github.com/ibndias/CIPHER accessed on 21 October 2024).